Expand the function f(z)=log 1+Z/ 1-Z in taylor series
Answer #1

ato 2 f(z) = log ( 1+ 2 ) = -log (2-2) fe) = -xoglı-2) - -|-2-2-2-2-2- >f(z) = (2+ z² + 3 + 24 + 25 3 4 5 2 your Ansuur. Get

Thank you

Please let me know if you need anything else from me

Upvote if helpful

Stay safe stay home

Have a great day